If Planet I is 31.5 million miles farther from the sun than Planet II, and Planet III is 24.1 million miles farther from the sun than Planet I. When the total of the distances for these three planets from the sun is 189.1 million miles, how far away from the sun is Planet III?
Answer by TimothyLamb(4379)

x = planet 1 distance from sun
y = planet 2 distance from sun
z = planet 3 distance from sun
---
x = y + 31.5
z = x + 24.1
x + y + z = 189.1
---
put the system of linear equations into standard form
---
x - y = 31.5
x - z = -24.1
x + y + z = 189.1
---
copy and paste the above standard form linear equations in to this solver:
https://sooeet.com/math/system-of-linear-equations-solver.php
---
solution:
x = planet 1 distance from sun = 65.5 million miles
y = planet 2 distance from sun = 34.0 million miles
z = planet 3 distance from sun = 89.6 million miles
---
Solve and graph linear equations:
https://sooeet.com/math/linear-equation-solver.php
---
Solve quadratic equations, quadratic formula:
https://sooeet.com/math/quadratic-formula-solver.php
---
Solve systems of linear equations up to 6-equations 6-variables:
https://sooeet.com/math/system-of-linear-equations-solver.php